OpenStudy (anonymous):

The density of solid Cr is 7.15 g/cm3. How many atoms are present per cubic centimeter of Cr

OpenStudy (anonymous):

\[1 cm^3 (\frac{7.15g Cr}{1 cm^3})(\frac{1 mole Cr}{51.996g Cr})(\frac{6.022x10^{23} atoms Cr}{1 mole Cr})=\]

OpenStudy (anonymous):

Perfect, Thanks a lot!